OpenStudy (anonymous):

every evening jenna empties her pocket and puts pockets and puts her change in a jar. at the end of the week she counts her money. one week she had 50 coins all of them were dimes and quarters. when she added them up she had a total of $10.25 let d=number of dimes and q=numbers of quarters. How many quarters did jenna have?

OpenStudy (anonymous):

help please

OpenStudy (campbell_st):

ok, you need to solve 2 equations simultaneously the coins d +q = 50 then for the value you have 0.1d + 0.25q = 10.25 take the 1st equation and rewrite it with d as the subject d = 50 - q substitute it into the 2nd equation 0.1(50 - q) + 0.25q = 10.25 distribute and collect like terms and then you will be able to solve for q when you know how many quarters, you will be able to work out the number of dimes as there are 50 coins in total. hope this helps